Question 224 of 417MediumMCQ
Two families, each with 2 children, boy/girl equally likely. Find the conditional probability all children are girls given at least two are girls.
Solutions · 1
Official solution
P(all 4 girls)=(21)4=161. P(at least 2 girls out of 4)=1−4C0(21)4−4C1(21)4=1611.
\[P = \frac{1/16}{11/16}=\frac{1}{11}\] Answer the question to unlock